# Staying on the Manifold: Geometry-Aware Noise Injection

### v1

@@ 3 Noise Injection Strategies
We consider three strategies of increasing complexity for geometry-aware input noise: tangential noise, geodesic noise and Brownian motion noise. The goal was to create noise injection techniques which either stay close to the manifold or, better, stay on the manifold.
@@ 3.1 Projected Tangent Space Noise
The simplest method we will try is Gaussian noise projected to the tangent space. Intuitively, this takes a noise sample in the ambient space, ${\boldsymbol{\epsilon}}\sim\mathcal{N}\left(\boldsymbol{0},\sigma^{2}\mathbb{I}_{D}\right)$, and pulls it closer to the manifold.
To isolate the tangential component $\boldsymbol{\epsilon}$$_{\top}$, we subtracting the component $\boldsymbol{\epsilon}_{\perp}$:
${\boldsymbol{\epsilon}}_{\top}={\boldsymbol{\epsilon}}-\sum_{i}\langle{\boldsymbol{\epsilon}},\boldsymbol{n}_{i}\rangle\cdot\boldsymbol{n}_{i},$
where $\{\boldsymbol{n}_{i}\}$ is a set of unit vectors spanning the normal space of $\mathcal{M}$. For more details we recommend the classic textbook [15]. Equivalently, the tangential noise can be defined with projection matrix, $\mathbf{P}=\mathbb{I}_{D}-\sum_{i}\boldsymbol{n}_{i}\boldsymbol{n}_{i}^{\top}$, as
$\boldsymbol{\epsilon}_{\top}=\mathbf{P}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}\left(\boldsymbol{0},\sigma^{2}\mathbb{I}_{D}\right).$
This allows for directly sampling tangent noise as
$\boldsymbol{\epsilon}_{\top}\sim\mathcal{N}\left(\boldsymbol{0},\sigma^{2}\mathbf{P}\right).$
@@
We now analyse how adding tangential noise $\boldsymbol{\epsilon}$$_{\top}$ affects the model $f_{{\boldsymbol{\theta}}}$. We proceed as in Subsection 2.1.
$\mathbb{E}[{\boldsymbol{\epsilon}}_{\top}^{\top}\mathbf{H}_{\mathcal{L}}$ ${\boldsymbol{\epsilon}}_{\top}]=\sum_{n=1}^{N}\mathbb{E}\left[{\boldsymbol{\epsilon}}_{\top}^{\top}\nabla_{x}f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n})\nabla_{x}f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n})^{\top}{\boldsymbol{\epsilon}}_{\top}\right]$
$+\frac{1}{2}\sum_{n=1}^{N}\mathbb{E}\left[{\boldsymbol{\epsilon}}_{\top}^{\top}\left(f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n})-y_{n}\right)\Delta_{x}f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n}){\boldsymbol{\epsilon}}_{\top}\right]$
$=\sum_{n=1}^{N}\lVert\nabla_{x}f_{{\boldsymbol{\theta}}}({\boldsymbol{x}})_{\top}\rVert^{2}.$
The second summand again vanishes if we assume that the model $f_{\boldsymbol{\theta}}$ interpolates the target values perfectly, that is, $f_{\boldsymbol{\theta}}({\boldsymbol{x}}_{n})=y_{n}$ for all $n=1,\dots,N$. When evaluating the first summand, we see that
@@ 3.2 Geodesic Noise
As explained in Subsection 2.2, at every $\boldsymbol{x}$$\in\mathcal{M}$, and for every ${\boldsymbol{v}}\in T_{{\boldsymbol{x}}}\mathcal{M}$ there exists a geodesic $\gamma:I\rightarrow\mathcal{M}$ such that $\gamma(0)={\boldsymbol{x}},$ and $\dot{\gamma}(0)={\boldsymbol{v}}.$ All manifolds in our paper are complete, and hence $I=\mathbb{R},$ and $\gamma$ can be extended to the whole of $\mathbb{R}.$ This allows us to generate points $\tilde{{\boldsymbol{x}}}$ near $\boldsymbol{x}$} by sampling initial velocities and mapping them to the manifold via the exponential map. We proceed as follows:
  1. 1.
    sample ${\boldsymbol{\epsilon}}\sim\mathcal{N}(0,\sigma^{2}\mathbb{I}_{D}),$
  2. 2.
    project $\boldsymbol{\epsilon}$} into the tangent space $T_{{\boldsymbol{x}}}\mathcal{M}$ as explained in Subsection 3.1, let
    ${\boldsymbol{\epsilon}}_{\top}=\mathbf{P}{\boldsymbol{\epsilon}},$
  3. 3.
    evaluate $\gamma$ at $|{\boldsymbol{\epsilon}}_{\top}$$|$, the new point is now
@@
For a small step size $\sigma$, we expect this to have a similar effect as the tangential noise but may improve robustness for increased step sizes. Details about the implementation of this method can be found in Appendix B.
@@ 3.3 Intrinsic Brownian Motion
Brownian motion (BM) is a stochastic process, which has first been used to describe the random movement of particles suspended in a fluid. Due to its occurrence in nature, this provides a realistic way of modelling how data points might move along a manifold. In the parameter space of a Riemannian manifold, the Brownian motion is defined by the following stochastic process [16]:
$du_{k}(t)$ $=\frac{1}{2}\frac{1}{\sqrt{\det g}}\sum_{l=1}^{d}\frac{\partial}{\partial u_{l}}\left(\sqrt{\det g}\cdot g^{kl}\right)dt$ (7)
$+\left(\sqrt{g^{-1}}dB(t)\right)_{k}.$
where $dB(t)$ is the standard Euclidean BM and $t$ is the time. The summands are referred to as the drift term and noise term, respectively. Since the Brownian motion on a manifold is generated by the Laplace-Beltrami operator [17], which is intrinsic, it is independent of the local charts [18].
@@ 3.4 Example: the Swiss roll
In this subsection, we will go through the computations for one example manifold, namely the Swiss roll. The derivations for the other manifolds follow a similarly approach. The Swiss roll is parametrised as follows:
$X:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3},\ \ X(u_{1},u_{2})=(au\sin{u_{1}},au_{1}\cos{u_{1}},u_{2}).$
Here, $a\in\mathbb{R}^{+}$ is a positive coefficient which determines how tightly the manifold is rolled. The metric is then given by
$\begin{pmatrix}a^{2}(1+u_{1}^{2})&0\\ 0&1\end{pmatrix}$.
@@
The unit normal vector at each point $(u,v)$ is given by
@@
To get tangential noise, we subtract the normal component from the Gaussian ambient noise, i.e.
${\boldsymbol{\epsilon}}_{\top}={\boldsymbol{\epsilon}}-\langle{\boldsymbol{\epsilon}},\boldsymbol{n}\rangle\cdot\boldsymbol{n}.$
@@
A curve on the manifold can be obtained by taking a curve $\alpha:I\rightarrow\mathbb{R}^{2}$ in $\mathbb{R}^{2}$ and mapping it on the manifold through $X.$ Let $\gamma:I\rightarrow\mathcal{M}$ be given by $\gamma(t)=X\circ\alpha(t).$
$\ddot{\alpha}_{1}(t)$ $=-\frac{\alpha_{1}(t)}{1+\alpha_{1}(t)^{2}}\dot{\alpha}_{1}(t)^{2},$
$\alpha_{2}(t)$ $=\alpha_{2}(0)+t\dot{\alpha}_{2}(0).$
@@
We start with a few pre-computations. It is clear that
$\det(g)$ $=a^{2}(1+u^{2}),$
$g^{-1}$ $=\begin{pmatrix}\frac{1}{a^{2}\left(1+u_{1}^{2}\right)}&0\\ 0&1\end{pmatrix}$.
Plugging it into Equation 7, we get:
$\begin{bmatrix}du_{1}\\ du_{2}\end{bmatrix}$ $=-\frac{dt}{2}\begin{pmatrix}\frac{u}{(1+u^{2})^{2}}\\ 0\end{pmatrix}+\sqrt{dt}\begin{pmatrix}\frac{1}{\sqrt{a^{2}\left(1+u^{2}\right)}}&0\\ 0&1\end{pmatrix}\tilde{{\boldsymbol{\epsilon}}}$.
@@
Figure 3: Brownian motion is computed from an initial point ( ) in the parameter space (left), then mapped to the manifold (right) via the chart $X$. The endpoint of the Brownian motion on the manifold ( ) acts as the noisy observation.
Figure 4: Illustration of the deformation process using a flow field $v_{t}$ for increasing time steps. The base point cloud is the sphere in $\mathbb{R}^{3}$.
Figure 5: Illustrations of the manifolds and functions on manifolds that we consider. From left: Sphere, SqueezedSphere, DeformedSphere, Bead, OnionRing and SwissRoll. Specifically, we use the deformation approach described in Section 4 to construct the DeformedSphere from a parametrised unit sphere in $\mathbb{R}^{3}$.
@@ 4 Deformation of a Manifold
In the following, we briefly elaborate on the underlying approach to constructing deformations of manifolds which we will use in Section 5.
We consider a vector field $v$ for defining a time-dependent diffeomorphism, $\phi:\mathcal{M}\times[0,T]\rightarrow\mathbb{R}^{D}$ that maps points from a parametrised manifold to a deformed version of the manifold, $\widetilde{\mathcal{M}}$. This is also known as a flow. The vector field induces the flow through an ordinary differential equation:
$\frac{d}{dt}\phi_{t}\left(\boldsymbol{x}\right)$ $=v_{t}\left(\phi_{t}\left({\boldsymbol{x}}\right)\right),$ (8)
$\phi_{0}\left(\boldsymbol{0}\right)$ $={\boldsymbol{x}},$
where $\boldsymbol{x}$$\in\mathcal{M}$ is a point on the parametrised manifold. As our manifold is defined through local coordinates, we can express points on the deformed manifold as
$\tilde{\boldsymbol{x}}=\phi_{T}\left(X\left(\boldsymbol{u}\right)\right)\in\widetilde{\mathcal{M}}$,
which is obtained by integrating the ODE (8) up to time $T$. We provide an illustration of such a deformation process for the sphere in Figure 4. The Jacobian of $\phi_{t}$ with respect to $u$ at $\boldsymbol{u}=X^{-1}({\boldsymbol{x}})$ is given by
@@
$\frac{d}{dt}\mathbf{J}_{\boldsymbol{u}}\left(t\right)$ $=\mathbf{J}_{v}\left(t\right)\mathbf{J}_{\boldsymbol{u}}\left(t\right),$
$\mathbf{J}_{\boldsymbol{u}}\left(0\right)$ $=\frac{\partial X\left(\boldsymbol{u}\right)}{\partial u},$
@@
$\tilde{g}=\mathbf{J}_{\boldsymbol{u}}\left(T\right)^{\top}\mathbf{J}_{\boldsymbol{u}}\left(T\right).$
We can now sample initial velocities on the tangent space $T_{\tilde{{\boldsymbol{x}}}}\widetilde{M}$ at $\tilde{{\boldsymbol{x}}}$, using (9) and generate geodesics on the deformed manifold $\widetilde{M}$.
This framework allows for highly expressive and flexible deformations of any parametrised manifold while ensuring invertibility. Previous research [19, 20] parametrise $v_{t,\boldsymbol{\theta}}$ with a neural network. Though we in practice only consider a fixed parametrisation of such a network to generate diffeomorphic deformations, our framework works for any map $v_{t}$. This opens new pathways to neural network settings where a learned flow approximates the underlying data manifold from which we can compute intrinsic geometric quantities, which we leave for future work.
@@
The Jacobian of the vector field, $v_{t}$, rarely has a closed form, however we can compute it efficiently using automatic differentiation. In practice, this allows us to evaluate derivatives of deformed manifolds with respect to the local coordinates of points on the manifold, without manually deriving the expressions. This algorithmic framework allows us to apply the technique to any manifold as long as some parametrisation is available and we have a differentiable ODE solver. In practice, we solve the flow equation numerically using an explicit Euler scheme and compute Jacobians and induced metrics with automatic differentiation. We remark that higher-order ODE solvers could be used for improved accuracy, yet the choice of the Euler scheme is based on practical challenges with existing toolboxes, for instance current incompatibility issues between existing libraries.
Table 1: Relative mean squared error to the baseline (B) model trained without adding noise. We report results for the optimal hyperparameter $\sigma^{2}$ for each strategy and manifold. We compare with ambient noise (A), tangent noise (T), geodesic noise (G) and Brownian motion noise (BM). We highlight the best strategy per manifold in bold. Adding noise does not improve performance for some manifolds, but results are included for completeness.

### v2

@@ 3 Noise Injection Strategies
We consider three strategies of increasing complexity for geometry-aware input noise: tangential noise, geodesic noise and Brownian motion noise. These noise injection strategies either stay close to the manifold or, better, stay on the manifold.
@@ 3.1 Projected Tangent Space Noise
One strategy is to project Gaussian noise to the tangent space. This takes a sample in the ambient space, ${\boldsymbol{\epsilon}}\sim\mathcal{N}\left(\boldsymbol{0},\sigma^{2}\mathbb{I}_{D}\right)$, and pulls it closer to the manifold. The tangential component $\boldsymbol{\epsilon}$$_{\top}$ is found by subtracting the orthogonal part, $\boldsymbol{\epsilon}$$_{\perp}$ from $\boldsymbol{\epsilon}$}:
${\boldsymbol{\epsilon}}_{\top}={\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\perp}={\boldsymbol{\epsilon}}-\sum_{i}\langle{\boldsymbol{\epsilon}},\boldsymbol{n}_{i}\rangle\cdot\boldsymbol{n}_{i}.$ (12)
Here, $\{\boldsymbol{n}_{i}\}$ is a set of unit vectors spanning the normal space of $\mathcal{M}$. For more details we recommend the classic textbook [absil]. Equivalently, the tangential noise can be defined as $\boldsymbol{\epsilon}_{\top}=\mathbf{P}\boldsymbol{\epsilon}$ with projection matrix $\mathbf{P}=\mathbb{I}_{D}-\sum_{i}\boldsymbol{n}_{i}\boldsymbol{n}_{i}^{\top}$. This allows for directly sampling tangent noise as $\boldsymbol{\epsilon}_{\top}\sim\mathcal{N}\left(\boldsymbol{0},\sigma^{2}\mathbf{P}\right).$
@@
We now analyse how adding tangential noise $\boldsymbol{\epsilon}$$_{\top}$ affects the model $f_{{\boldsymbol{\theta}}}$. We proceed as in Subsection 2.1 and observe that
$\mathbb{E}[{\boldsymbol{\epsilon}}_{\top}^{\top}\mathbf{H}_{\mathcal{L}}{\boldsymbol{\epsilon}}_{\top}]=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}[{\boldsymbol{\epsilon}}_{\top}^{\top}\nabla_{x}f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n})\nabla_{x}f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n})^{\top}{\boldsymbol{\epsilon}}_{\top}]$
$+\frac{1}{2N}\sum_{n=1}^{N}\mathbb{E}[{\boldsymbol{\epsilon}}_{\top}^{\top}\left(f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n})-{\boldsymbol{y}}_{n}\right)\Delta_{x}f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n}){\boldsymbol{\epsilon}}_{\top}]$ (13)
The second summand again vanishes if we assume that the model $f_{\boldsymbol{\theta}}$ interpolates the target values perfectly, that is, $f_{\boldsymbol{\theta}}({\boldsymbol{x}}_{n})={\boldsymbol{y}}_{n}$ for all $n=1,\dots,N$. When evaluating the first summand, we use an orthogonal decomposition of the gradient to see that
@@ 3.2 Geodesic Noise
As explained in Subsection 2.2, at every $\boldsymbol{x}$$\in\mathcal{M}$, and for every ${\boldsymbol{v}}\in T_{{\boldsymbol{x}}}\mathcal{M}$ there exists a geodesic $\gamma:I\rightarrow\mathcal{M}$ such that $\gamma(0)={\boldsymbol{x}},$ and $\dot{\gamma}(0)={\boldsymbol{v}}.$ All manifolds in our paper are complete, and hence $I=\mathbb{R},$ and $\gamma$ can be extended to the whole of $\mathbb{R}.$ This allows us to generate points $\tilde{{\boldsymbol{x}}}$ near $\boldsymbol{x}$} by sampling initial velocities and mapping them to the manifold via the exponential map. We proceed as follows: first, sample a velocity $\boldsymbol{\epsilon}$$_{\top}$ in the tangent space $T_{{\boldsymbol{x}}}\mathcal{M}$ as explained in Subsection 3.1, next, evaluate $\gamma$ at ${\boldsymbol{\epsilon}}_{\top}||$ to get the geodesic noise sample,
@@
For a small step size $\sigma$, we expect this to have a similar effect as the tangential noise but may improve robustness for increased step sizes. Details about the implementation can be found in Appendix A.
@@ 3.3 Intrinsic Brownian Motion
Brownian motion is a stochastic process, which has been used to describe random movement of particles suspended in a fluid. Due to its occurrence in nature, this provides a realistic way of modelling how data points might move on a manifold. In the parameter space of a Riemannian manifold, Brownian motion is defined by the following stochastic process [hsu2008brief]:
$du_{k}(t)$ $=\frac{1}{2}\frac{1}{2\sqrt{\det g}}\sum_{l=1}^{d}\frac{\partial}{\partial u_{l}}\left(\sqrt{\det g}\cdot g^{kl}\right)dt$
$+\left(\sqrt{g^{-1}}dB(t)\right)_{k}$ (17)
where $dB(t)$ is Euclidean Brownian motion and $t$ is the time. The summands are referred to as the drift and noise term, respectively. Since Brownian motion on a manifold is generated by the Laplace-Beltrami operator [hsu2002stochastic], which is intrinsic, it is independent of the chart [ikeda2014stochastic]. We visualise the strategy in Figure 3.
@@ 3.4 Example: the Swiss Roll
We will now do the computations for one example manifold, namely the Swiss roll. This manifold is parameterised by a chart $X:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$ as follows:
$X\left(u_{1},u_{2}\right)=\left(au_{1}\sin{u_{1}},au_{1}\cos{u_{1}},u_{2}\right).$
Here, $a\in\mathbb{R}^{+}$ is a coefficient which determines how tightly the manifold is rolled. The metric $g$ is then
$g=\operatorname{diag}\begin{pmatrix}a^{2}\left(1+u_{1}^{2}\right),1\end{pmatrix}$.
@@
The unit normal vector at each point $X(u_{1},u_{2})$ is given by
@@
Following Subsection 3.1, we generate tangential noise from the normal vector and a Gaussian sample.
@@
A curve on the manifold $\gamma:I\rightarrow\mathcal{M}$ can be obtained by taking a curve $\alpha:I\rightarrow\mathbb{R}^{2}$ in the parameter space $\mathbb{R}^{2}$ and mapping it on the manifold via $X.$ For the Swiss roll, the Geodesic Equation, i.e. Equation 9, is equivalent to
$\ddot{\alpha}_{1}(t)=-\frac{\alpha_{1}(t)\dot{\alpha}_{1}(t)^{2}}{1+\alpha_{1}(t)^{2}},\quad\ddot{\alpha}_{2}(t)=\alpha_{2}(0)+t\dot{\alpha}_{2}(0).$
@@
For the metric $g$, we have
$\det(g)=a^{2}(1+u_{1}^{2})\quad\text{and}\quad g^{-1}=\operatorname{diag}\begin{pmatrix}\frac{1}{a^{2}(1+u_{1}^{2})},1\end{pmatrix}$.
Plugging these quantities into Equation 3.3, we get:
$\begin{bmatrix}du_{1}\\ du_{2}\end{bmatrix}$ $=-\frac{dt}{2}\begin{bmatrix}\frac{u_{1}}{(1+u_{1}^{2})^{2}}\\ 0\end{bmatrix}+\sqrt{dt}\begin{bmatrix}\frac{1}{\sqrt{a^{2}\left(1+u_{1}^{2}\right)}}\\ 1\end{bmatrix}\odot\tilde{{\boldsymbol{\epsilon}}}$.
@@
Figure 3: Brownian motion from an initial point ( ) in the parameter space (left), and mapped to the manifold (right) via the chart $X$. The endpoint of the Brownian motion on the manifold ( ) acts as the noisy observation.
@@ 4 Deformation of a Manifold
We briefly elaborate on an approach to deform parameterised manifolds, which we use in Section 5. We consider a vector field $v$ for defining a time-dependent diffeomorphism, $\phi:\mathcal{M}\times[0,T]\rightarrow\mathbb{R}^{D}$ that maps points from a parameterised manifold to a deformed version of the manifold, $\widetilde{\mathcal{M}}$. This is also known as a flow. The vector field $v$ induces the flow through an ordinary differential equation:
$\phi_{0}\left(\boldsymbol{0}\right)={\boldsymbol{x}},\quad\frac{d}{dt}\phi_{t}\left(\boldsymbol{x}\right)=v_{t}\left(\phi_{t}\left({\boldsymbol{x}}\right)\right),$ (18)
where $\boldsymbol{x}$$\in\mathcal{M}$ is a point on the parameterised manifold. We can then express points on the deformed manifold through the local coordinates of the parameterised manifold as $\tilde{\boldsymbol{x}}=\phi_{T}\left(X\left(\boldsymbol{u}\right)\right)\in\widetilde{\mathcal{M}}$, which is obtained by integrating the ODE up to time $T$. We provide an illustration of such a deformation process for the sphere in Figure 4. The Jacobian of $\phi_{t}$ with respect to $u$ at $\boldsymbol{u}=X^{-1}({\boldsymbol{x}})$ is given by
@@
$\mathbf{J}_{\boldsymbol{u}}\left(0\right)=\frac{\partial X\left(\boldsymbol{u}\right)}{\partial u},\quad\frac{d}{dt}\mathbf{J}_{\boldsymbol{u}}\left(t\right)=\mathbf{J}_{v}\left(t\right)\mathbf{J}_{\boldsymbol{u}}\left(t\right),$ (19)
@@
$\tilde{g}=\mathbf{J}_{\boldsymbol{u}}\left(T\right)^{\top}\mathbf{J}_{\boldsymbol{u}}\left(T\right).$ (20)
This allows sampling vectors on the tangent space $T_{\tilde{{\boldsymbol{x}}}}\widetilde{M}$ at $\tilde{{\boldsymbol{x}}}$ and generating geodesics or Brownian motion on the deformed manifold $\widetilde{M}$ by pulling the metric back to the parameter space. This framework allows for highly expressive and flexible deformations of any parameterised manifold while ensuring invertibility. Previous research [chen2018neural, lipman2022flow] parameterise $v_{t,\boldsymbol{\theta}}$ with a neural network. Though we in practice only consider a fixed parameterisation of such a network, our framework works for any map $v_{t}$. This opens new pathways to neural network settings where a learnt flow approximates the data manifold from which we can then compute intrinsic geometric quantities, which we leave for future work.
Figure 4: The deformation process of the sphere in $\mathbb{R}^{3}$ for increasing time steps of using a flow field $v_{t}$.
@@
The Jacobian of the vector field, $v_{t}$, rarely has a closed form, however we can compute it efficiently using automatic differentiation (AD) with e.g. JAX or PyTorch. In practice, this allows us to evaluate derivatives of deformed manifolds with respect to the local coordinates of points on the manifold, without manually deriving the expressions. This algorithmic framework allows us to apply the technique to any manifold as long as some parameterisation is available and we have a differentiable ODE solver. In practice, we solve the flow equation numerically using an Euler scheme and compute Jacobians and induced metrics with AD. We remark that higher-order ODE solvers can be used for improved accuracy, yet the Euler scheme was chosen due to challenges with current toolboxes, specifically incompatibility issues between libraries.
Table 1: Mean squared error relative to the baseline (B) model trained without adding noise. We report results for the optimal hyperparameter $\sigma^{2}$ for each strategy and manifold. We compare with ambient noise (A), tangent noise (T), geodesic noise (G) and Brownian motion noise (BM). We highlight the best strategy per manifold in bold. Adding noise does not improve performance for some manifolds, but results are included for completeness. We include illustrations of the manifolds and functions on manifolds that we consider. The deformation approach described in Section 4 is used to construct the DeformedSphere from a parameterised unit sphere in $\mathbb{R}^{3}$.
@@
Manifold
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]